# Dual-Flow: Transferable Multi-Target, Instance-Agnostic Attacks via In-the-wild Cascading Flow Optimization

### v1

@@ Abstract
Adversarial attacks are widely used to evaluate model robustness, and in black-box scenarios, the transferability of these attacks becomes crucial. Existing generator-based attacks have excellent generalization and transferability due to their instance-agnostic nature. However, when training generators for multi-target tasks, the success rate of transfer attacks is relatively low due to the limitations of the model’s capacity. To address these challenges, we propose a novel Dual-Flow framework for multi-target instance-agnostic adversarial attacks, utilizing Cascading Distribution Shift Training to develop an adversarial velocity function. Extensive experiments demonstrate that Dual-Flow significantly improves transferability over previous multi-target generative attacks. For example, it increases the success rate from Inception-v3 to ResNet-152 by 34.58%. Furthermore, our attack method, such as adversarially trained models, shows substantially stronger robustness against defense mechanisms.
@@ 1 Introduction
Adversarial attacks can be classified into instance-specific and instance-agnostic approaches. Instance-specific attacks [10, 57, 14] optimize perturbations for each input using surrogate model gradients but often suffer from poor transferability. Instance-agnostic attacks [55, 38, 59, 15] generalize perturbations over the dataset, leading to stronger black-box transferability. These methods typically rely on universal adversarial perturbations [36, 60] or generative models [40, 37].
@@
Diffusion models [25, 45, 42] offer strong generative capabilities, making them a promising tool for adversarial attacks. However, existing diffusion-based attacks rely on iterative optimization using the victim model’s gradients, preventing them from being truly instance-agnostic. Moreover, choosing between stochastic and deterministic sampling methods for adversarial perturbation generation remains an open challenge.
@@ 2.1 Instance-Agnostic Attacks
Consider a white-box image classifier characterized by the parameter $\theta$, denoted as $f:\mathcal{X}\to\mathcal{Y}$, where the input space $\mathcal{X}\subset\mathbb{R}^{C\times H\times W}$ corresponds to the image domain, and the output space $\mathcal{Y}\subset\mathbb{R}^{L}$ represents the confidence scores across various classes. Here, $L$ denotes the total number of classes. Given an original image $\mathbf{x}\in\mathcal{X}$ and a target class $c\in\mathcal{C}$, the goal of transferable multi-target generative attack is to generate the perturbation $\boldsymbol{\delta}=G(\mathbf{x},c)$ and the pertubed image $\boldsymbol{\mathbf{x^{\epsilon}}}=\mathbf{x}+\boldsymbol{\delta}$, in such a way that an unseen victim classifier $F$ predicts $c$ for the perturbed image, i.e., $\arg\max_{i\in\mathcal{C}}F(\mathbf{x}^{\epsilon})_{i}=c$. Here $G$ is the generator trained on the known model $f_{\theta}$. To ensure that the manipulated images remain visually indistinguishable from the originals, the perturbation is constrained using the $l_{\infty}$ norm such that $\|\boldsymbol{\mathbf{x}-\mathbf{x}^{\epsilon}}\|_{\infty}=\|\boldsymbol{\delta}\|_{\infty}<\epsilon$.
@@ 2.2 Diffusion Models and Flow-based Models
Diffusion Mmodels in Adversarial Attack.
Currently, diffusion models have found some applications in adversarial attacks. Some utilize diffusion models to create unrestricted adversarial examples[5, 6], while others perform instance-specific attacks[58, 4]. However, they all rely on iterative optimization using the target model’s gradients while generating adversarial examples, thus not qualifying as instance-agnostic attacks.
@@ 3 Dual-Flow for Adversarial Attack
We propose a Dual-Flow pipeline designed to transform an image $\mathbf{x}\in\mathcal{X}$ through a perturbed distribution $\mathcal{X}_{\tau}$ and ultimately into a constrained output space $\mathcal{X}^{\epsilon}$, which is $\{\mathbf{x}^{\epsilon}|\exists\mathbf{x}\in\mathcal{X},\|\mathbf{x^{\epsilon}-\mathbf{x}}\|_{\infty}<\epsilon\}$. By construction, $\mathcal{X}^{\epsilon}$ enforces a maximum $\ell_{\infty}$ perturbation of $\epsilon$. Unlike approaches that rely on guidance from a victim model, our inference process is entirely model-agnostic.
@@ 3.1 A Construction of Better Extend Flow
$\begin{split}j=-\mathbb{CE}(f(\mathbf{x})),c),\end{split}$ (3)
where $f$ is the classifier and $c$ is the target label.
Proposition 3.1 (Morse Flow Construction).
Under maid assumptions on the $\mathcal{X}^{\epsilon}$ and function $j$, There exits $\epsilon>0$, a unique smooth flow
@@
    $j(\Phi(\mathbf{x},\epsilon))\geq j(\mathbf{x})$ for all $x\in\mathcal{X}^{\epsilon}$, and $>$ holds almost everywhere if $j$ is not trivial
@@
A detailed proof is provided in Appendix A.1. Proposition 3.1 indicates we can find better extend flow to hack $j$ by $\nabla j$. In the following paragraphs, we will construct a concrete algorithm to realize it along with an existing flow.
@@ 3.2 Cascading Distribution Shift Training
  Input: $\tau=N\delta$, stepsize $\delta$, model param. $\phi$, $\theta$, victim model $f$, target labels set $\mathcal{C}$, training dataset $\{\mathbf{I}^{i}\}_{i\in\mathcal{I}}$, learning rate $l_{r}$
@@
           $\widehat{\mathbf{x}_{0}}^{i,j,k}=\operatorname{clip}\left(\widehat{\mathbf{x}_{0}}^{i,j,k},\mathbf{x}^{i,j,k}-\epsilon,\mathbf{x}^{i,j,k}+\epsilon\right)$
@@
Although the in-the-wild ODE trajectory is deterministic, obtaining exact intermediate samples remains a significant challenge, which poses difficulties for our approach. To address this issue, we propose the Cascading Distribution Shift Training Algorithm, specifically designed to enhance adversarial attack efficacy through two key mechanisms: (1) Enforcing a cascading hacking effect, where each perturbation step incrementally contributes to misleading the victim model. (2) Ensuring the final perturbation follows the prescribed $\ell_{\infty}$ constraint. Further details are provided in Algorithm 1, where $l$ denotes the conditional input in real diffusion models.
Our proposed training framework not only circumvents the challenge of inaccessible intermediate timesteps but also offers additional advantages. Proposition 3.2 formalizes that our algorithm progressively refines a coarse-to-fine representation, thus effectively leveraging information from ahead timesteps. More concretely, due to the continuity of the ODE, our training algorithm enables a cascading optimization within an increasingly refined space.
Proposition 3.2 (Cascading Improvement at Adjoint Timesteps).
@@ 3.3 Dual-Flow Sampling
  Input: $\tau=N\delta$, stepsize $\delta$, image $\mathbf{I}$, target label $c$ Dual-Flow $\{\mathbf{v}_{\phi},\mathbf{v}_{\theta}\}$
@@
  $\mathbf{x}_{0}^{i,j,k}=\operatorname{clip}\left(\mathbf{x}_{0}^{i,j,k},\mathbf{x}^{i,j,k}-\epsilon,\mathbf{x}^{i,j,k}+\epsilon\right)$
@@
First, Cascading SDE introduces a random term, which may make it more difficult to construct the Cascading Improvement relationship as stated in Proposition 3.2 for Cascading ODE. Second, sampling with SDEs tends to produce unstable results, where larger step sizes exacerbate accumulated errors, further impacting reliability, as demonstrated in our experiments.
@@ 4 Experiments
[…]
@@ 4.1 Experimental Settings
[…]
@@
Following previous work [59, 16, 15], we choose Res-152 and Inc-v3 as substitute models to train our model. The perturbation budget $\epsilon$ is 16/255. We conduct 50k steps of training for multi-target tasks. To compare our method with other single-target attacks, we further fine-tune our model for an additional 10k steps to specialize in a single target class(more details provided in Appendix C.2). For multi-target training, we use a learning rate of $52.5e-$ 5 and a batch size of 8. For single-target fine-tuning, the learning rate is set to $51e-$ 5 with a batch size of 4.
@@ 4.2 Transferability Evaluation
We initially conduct attacks on normally trained models to evaluate the performance of multi-target attacks. The results in Table 1 show that our proposed Dual-Flow method exhibits significantly superior transferability, outperforming state-of-the-art instance-specific and instance-agnostic methods. Specifically, our method achieves an average ASR improvement of 21.16% and 12.36% over CGNC [15] using Inc-v3 and Res-152 as surrogate models, respectively, on black-box models. Notably, instance-specific methods, despite higher success rates in white-box settings, tend to overfit the surrogate models’ classification boundaries, resulting in poor performance when transferred to black-box models.
@@
Figure 2: Visualization results of different input images targeting various classes. For each text prompt of the target class, the left column displays the adversarial examples generated before clipping, the middle column shows the adversarial examples after clipping, and the right column presents the corresponding adversarial perturbations, which represent the differences between the clipped adversarial examples and the original images. Note that the perturbations are scaled to a range between 0 and 1. The surrogate model used is Inc-v3.
@@ 4.3 Attack Under Defense Strategies
Figure 3: A comparison of CGNC and our method regarding attack success rates against various input processing defense strategies. The results against JPEG compression are shown in (a), while (b) presents the outcomes against different input smoothing methods. Inc-v3 is used as the surrogate model in this evaluation, and Inc-Res-v2, along with DN-121, are the target models.
Robustly Trained Networks.
We first consider attacking six robustly trained networks, with results in Table 3. Attacking robustness-augmented models is challenging, as previous methods see a significant drop in success rates. However, our approach consistently misleads black-box classifiers into predicting the specified classes, showing marked improvement over earlier multi-target methods. Notably, using Inc-v3 as the surrogate model, the average attack success rate against the six robust models increases significantly from 28.60% to 62.28%, highlighting our method’s effectiveness.
@@
We compared our method’s performance with the state-of-the-art multi-target attack method CGNC against input preprocessing defenses, such as image smoothing [9] and JPEG compression [12]. As shown in Figure 3, our method consistently outperforms CGNC under these defenses. For example, using Inc-v3 as the surrogate model and DN-121 as the target model, our method achieves a 52.99% success rate under Gaussian smoothing, compared to CGNC’s 28.27%. This highlights the superior effectiveness of our approach in overcoming input preprocessing defenses.
@@ 4.4 Visualization
To gain a deeper understanding of the effectiveness of our method, we visualized both the unclipped and clipped samples generated by our approach. Additionally, for consistency with other perturbation-based attack methods, we visualized the equivalent adversarial perturbations, defined as the pixel differences between the clipped samples and the clean images. As illustrated in Figure 2, our method first transforms the original image into one that maintains a similar layout and color scheme but becomes semantically closer to the target class. This transformed image is subsequently clipped to ensure its pixel differences from the original image remain within the epsilon bound. Visually, it is evident that the clipped image retains substantial semantic features of the target class. Notably, the adversarial perturbations also exhibit distinct semantic patterns aligned with the target class, further validating the effectiveness of our approach.
@@ 4.5 Ablation Study
Table 4: The multi-target attack success rates of several variants of our method. The surrogate model used is Res-152. The white-box attack success rate refers to the performance on Res-152, while the black-box attack success rate represents the average performance across six black-box models.
@@
Figure 4: The multi-target black-box attack success rates of several variants of our method. The surrogate model used is Res-152.
5 Discussion
Potential Advantage of SDE.
@@ 6 Conclusion
We have introduced Dual-Flow, a novel framework for highly transferable multi-target adversarial attacks. By employing Cascading Distribution Shift Training to develop an adversarial velocity function, our approach addresses the limitations of existing methods. Extensive experimental results demonstrate that Dual-Flow achieves remarkable improvements in transferability and robustness compared to previous multi-target generative attacks. These findings highlight the potential of Dual-Flow as a powerful tool for evaluating and improving the robustness of machine learning models.
Acknowledgement

### v3

@@ Abstract
Adversarial attacks are widely used to evaluate model robustness, and in black-box scenarios, the transferability of these attacks becomes crucial. Existing generator-based attacks have excellent generalization and transferability due to their instance-agnostic nature. However, when training generators for multi-target tasks, the success rate of transfer attacks is relatively low due to the limitations of the model’s capacity. To address these challenges, we propose a novel Dual-Flow framework for multi-target instance-agnostic adversarial attacks, utilizing Cascading Distribution Shift Training to develop an adversarial velocity function. Extensive experiments demonstrate that Dual-Flow significantly improves transferability over previous multi-target generative attacks. For example, it increases the success rate from Inception-v3 to ResNet-152 by 34.58%. Furthermore, our attack method shows substantially stronger robustness against defense mechanisms, such as adversarially trained models. The code of Dual-Flow is available at: https://github.com/Chyxx/Dual-Flow.
@@ 1 Introduction
Adversarial attacks can be classified into instance-specific and instance-agnostic approaches. Instance-specific attacks [12, 67, 16] optimize perturbations for each input image using victim model gradients but often suffer from poor transferability. Instance-agnostic attacks [65, 42, 69, 17] generalize perturbations over the dataset, leading to stronger black-box transferability. These methods typically rely on universal adversarial perturbations [41, 70] or generative models [46, 43].
@@
Diffusion models [29, 51, 48] offer strong generative capabilities, making them a promising tool for adversarial attacks. However, current diffusion-based attacks are all instance-specific methods, which means they need to access the target model’s gradient information during inference. Moreover, choosing between stochastic and deterministic sampling methods for adversarial perturbation generation remains an open challenge.
@@ 2.1 Instance-Agnostic Attacks
Consider a white-box image classifier characterized by the parameter $\theta$, denoted as $f:\mathcal{X}\to\mathcal{Y}$, where the input space $\mathcal{X}\subset\mathbb{R}^{C\times H\times W}$ corresponds to the image domain, and the output space $\mathcal{Y}\subset\mathbb{R}^{L}$ represents the confidence scores across various classes. Here, $L$ denotes the total number of classes. Given an original image $\mathbf{x}\in\mathcal{X}$ and a target class $c\in\mathcal{C}$, the goal of transferable multi-target generative attack is to generate the perturbation $\bm{\delta}=G(\mathbf{x},c)$ and the pertubed image $\bm{\mathbf{x^{\epsilon}}}=\mathbf{x}+\bm{\delta}$, in such a way that an unseen victim model $F$ predicts $c$ for the perturbed image, i.e., $\arg\max_{i\in\mathcal{C}}F(\mathbf{x}^{\epsilon})_{i}=c$. Here $G$ is the generator trained on the known source model $f_{\theta}$. To ensure that the manipulated images remain visually indistinguishable from the originals, the perturbation is constrained using the $l_{\infty}$ norm such that $\|\bm{\mathbf{x}-\mathbf{x}^{\epsilon}}\|_{\infty}=\|\bm{\delta}\|_{\infty}<\epsilon$.
@@ 2.2 Diffusion Models and Flow-based Models
Diffusion Models in Adversarial Attack.
Currently, diffusion models have found some applications in adversarial attacks. Some utilize diffusion models to create unrestricted adversarial examples[5, 8], while others perform instance-specific attacks[68, 4]. However, they all rely on iterative optimization using the classifier model’s gradients while generating adversarial examples, thus not qualifying as instance-agnostic attacks.
@@ 3 Dual-Flow for Adversarial Attack
We propose a Dual-Flow pipeline designed to transform an image $\mathbf{x}\in\mathcal{X}$ through a perturbed distribution $\mathcal{X}_{\tau}$ and ultimately into a constrained output space $\mathcal{X}^{\epsilon}$, which is $\{\mathbf{x}^{\epsilon}|\exists\mathbf{x}\in\mathcal{X},\|\mathbf{x^{\epsilon}-\mathbf{x}}\|_{\infty}<\epsilon\}$. By construction, $\mathcal{X}^{\epsilon}$ enforces a maximum $\ell_{\infty}$ perturbation of $\epsilon$. Our method is instance-agnostic, which means that once training is completed, our model can generate adversarial examples during inference without requiring any access to the classifier model.
@@ 3.1 A Construction of Better Extend Flow
$\begin{split}j=-\mathbb{CE}(f(\mathbf{x}),c),\end{split}$ (3)
where $f$ is the source model and $c$ is the target label.
Proposition 1 (Morse Flow Construction).
Under mild assumptions on the $\mathcal{X}^{\epsilon}$ and the function $j$, there exists $\epsilon>0$, a unique smooth flow.
@@
    $j(\Phi(\mathbf{x},\epsilon))\geq j(\mathbf{x})$ for all $x\in\mathcal{X}^{\epsilon}$, and >> holds almost everywhere if $j$ is not trivial
@@
A detailed proof is provided in Appendix A.1. Proposition 1 indicates we can find better extend flow to hack $j$ by $\nabla j$. In the following paragraphs, we will construct a concrete algorithm to realize it along with an existing flow.
@@ 3.2 Cascading Distribution Shift Training
Although the in-the-wild ODE trajectory is deterministic, obtaining exact intermediate samples remains a significant challenge, which poses difficulties for our approach. To address this issue, we propose the Cascading Distribution Shift Training Algorithm, specifically designed to enhance adversarial attack efficacy through two key mechanisms: (1) Enforcing a cascading hacking effect, where each perturbation step incrementally contributes to misleading the source model. (2) Ensuring the final perturbation follows the prescribed $\ell_{\infty}$ constraint. Further details are provided in Algorithm 1, where $l$ denotes the conditional input in real diffusion models.
@@
Input: $\tau=N\delta$, stepsize $\delta$, model param. $\phi$, $\theta$, source model $f$, target labels set $\mathcal{C}$, training dataset $\{\mathbf{I}^{i}\}_{i\in\mathcal{I}}$, learning rate $l_{r}$
@@
    $\widehat{\mathbf{x}_{0}}=\operatorname{clip}\left(\widehat{\mathbf{x}_{0}},\mathbf{x}-\epsilon,\mathbf{x}+\epsilon\right)$
@@
Our proposed training framework not only circumvents the challenge of inaccessible intermediate timesteps but also offers additional advantages. Proposition 2 formalizes that our algorithm progressively refines a coarse-to-fine representation, thus effectively leveraging information from ahead timesteps. More concretely, due to the continuity of the ODE, our training algorithm enables a cascading optimization within an increasingly refined space.
Theorem 2 (Cascading Improvement at Adjoint Timesteps).
@@ 3.3 Dual-Flow Sampling
Input: $\tau=N\delta$, stepsize $\delta$, image $\mathbf{I}$, target label $c$, Dual-Flow $\{\mathbf{v}_{\phi},\mathbf{v}_{\theta}\}$
@@
$\mathbf{x}_{0}=\operatorname{clip}\left(\mathbf{x}_{0},\mathbf{x}-\epsilon,\mathbf{x}+\epsilon\right)$
@@
First, Cascading SDE introduces a random term, which may make it more difficult to construct the Cascading Improvement relationship as stated in Proposition 2 for Cascading ODE. Second, sampling with SDEs tends to produce unstable results, where larger step sizes exacerbate accumulated errors, further impacting reliability, as demonstrated in our experiments.
@@ 4 Experiments
Source Method $\textrm{Inc-v3}_{\textrm{adv}}$ $\textrm{IR-v2}_{\textrm{ens}}$ $\textrm{Res50}_{\textrm{SIN}}$ $\textrm{Res50}_{\textrm{IN}}$ $\textrm{Res50}_{\textrm{fine}}$ $\textrm{Res50}_{\textrm{Aug}}$ Average
@@ 4.1 Experimental Settings
For further evaluation, we also analyze the performance of our method on robustly trained models. These include adv-Inception-v3 ($\textrm{Inc-v3}_{\textrm{adv}}$) [22], ens-adv-Inception-ResNet-v2 ($\textrm{IR-v2}_{\textrm{ens}}$) [26], and several robustly trained ResNet-50 models. The ResNet-50 variants are: $\textrm{Res50}_{\textrm{SIN}}$ (trained on stylized ImageNet), $\textrm{Res50}_{\textrm{IN}}$ (trained on a mixture of stylized and Nature ImageNet), $\textrm{Res50}_{\textrm{fine}}$ (further fine-tuned with an auxiliary dataset [21]), and $\textrm{Res50}_{\textrm{Aug}}$ (trained with advanced data augmentation techniques from Augmix [28]).
@@
Following previous work [69, 18, 17], we choose Res-152 and Inc-v3 as source models to train our model. The perturbation budget $\epsilon$ is 16/255. We conduct 50k steps of training for multi-target tasks. To compare our method with other single-target attacks, we further fine-tune our model for an additional 10k steps to specialize in a single target class(more details provided in Appendix D.2). For multi-target training, we use a learning rate of $2.5\times 10^{-5}$ and a total batch size of 8 (distributed across two NVIDIA RTX 3090 GPUs, each with 24GB memory and batch size 4). Training under this setting takes approximately one day to complete. For single-target fine-tuning, we set the learning rate to $1\times 10^{-5}$ and a batch size of 4, conducted on a single NVIDIA RTX 3090 GPU, which requires approximately 4 hours. In total, the experiments involve approximately 160 GPU hours.
@@ 4.2 Transferability Evaluation
We initially conduct attacks on normally trained models to evaluate the performance of multi-target attacks. The results in Table 1 show that our proposed Dual-Flow method exhibits significantly superior transferability, outperforming state-of-the-art instance-specific and instance-agnostic methods. Specifically, our method achieves an average ASR improvement of 21.16% and 12.36% over CGNC [17] using Inc-v3 and Res-152 as source models, respectively, on black-box models. Notably, instance-specific methods, despite higher success rates in white-box settings, tend to overfit the source models’ classification boundaries, resulting in poor performance when transferred to black-box models.
@@
Figure 2: Visualization results of different input images targeting various classes. For each text prompt of the target class, the left column displays the adversarial examples generated before clipping, the middle column shows the adversarial examples after clipping, and the right column presents the corresponding adversarial perturbations, which represent the differences between the clipped adversarial examples and the original images. Note that the perturbations are scaled to a range between 0 and 1. The source model used is Inc-v3.
@@ 4.3 Attack Under Defense Strategies
Robustly Trained Networks.
We first consider attacking six robustly trained networks, with results in Table 3. Attacking robustness-augmented models is challenging, as previous methods see a significant drop in success rates. However, our approach consistently misleads black-box classifiers into predicting the specified classes, showing marked improvement over earlier multi-target methods. Notably, using Inc-v3 as the source model, the average attack success rate against the six robust models increases significantly from 28.60% to 62.28%, highlighting our method’s effectiveness.
@@
Figure 3: A comparison of CGNC and our method regarding attack success rates against various input processing defense strategies. The results against JPEG compression are shown in (a), while (b) presents the outcomes against different input smoothing methods. Inc-v3 is the source model and Inc-Res-v2, along with DN-121, are the target models.
@@
We compared our method’s performance with the state-of-the-art multi-target attack method CGNC against input preprocessing defenses, such as image smoothing [11] and JPEG compression [14]. As shown in Figure 3, our method consistently outperforms CGNC under these defenses. For example, using Inc-v3 as the source model and DN-121 as the target model, our method achieves a 52.99% success rate under Gaussian smoothing, compared to CGNC’s 28.27%. This highlights the superior effectiveness of our approach in overcoming input preprocessing defenses.
@@ 4.4 Visualization
To gain a deeper understanding of the effectiveness of our method, we visualized both the unclipped and clipped samples generated by our approach. Additionally, for consistency with other perturbation-based attack methods, we visualized the equivalent adversarial perturbations, defined as the pixel differences between the clipped samples and the clean images. As illustrated in Figure 2, our method first transforms the original image into one that maintains a similar layout and color scheme but becomes semantically closer to the target class. This transformed image is subsequently clipped to ensure its pixel differences from the original image remain within the epsilon bound. Visually, the clipped image retains substantial semantic features of the target class. Notably, the adversarial perturbations also exhibit distinct semantic patterns aligned with the target class, further validating the effectiveness of our approach.
@@ 4.5 Ablation Study
Figure 4: The multi-target black-box attack success rates of several variants of our method. The source model used is Res-152.
5 Discussion
Potential Advantage of SDE.
Table 4: The multi-target attack success rates of several variants of our method. The source model used is Res-152. The white-box attack success rate refers to the performance on Res-152, while the black-box attack success rate represents the average performance across six black-box models.
@@ 6 Conclusion
We have introduced Dual-Flow, a novel framework for highly transferable multi-target adversarial attacks. By employing Cascading Distribution Shift Training to develop an adversarial velocity function, our approach addresses the limitations of existing methods. Extensive experimental results demonstrate that Dual-Flow achieves remarkable improvements in transferability and robustness compared to previous multi-target generative attacks. These findings highlight the potential of Dual-Flow as a powerful tool for evaluating and improving the robustness of generative models.
Acknowledgments and Disclosure of Funding